Directions: Provide the answer to the following financial ratios problem:

NOTE: USE UP TO TWO DECIMAL PLACES ONLY


6. GROSS PROFIT MARGIN= 0.60

  INVENTORY TURNOVER= 3

  AVERAGE INVENTORY= P 450,000

  NET PROFIT MARGIN= 0.12

  NET PROFIT= ?


7. CURRENT RATIO = 4

  DEBT RATIO= 0.2

  EQUITY= P 600,000

  NONCURRENT ASSETS= P 200,000

  NONCURRENT LIABILITIES= ?


8. AVERAGE DAYS IN INVENTORY= 20

  INVENTORY TURNOVER= ?


9. CURRENT LIABILITIES= 50% OF TOTAL ASSETS

  CURRENT ASSETS= 75% OF TOTAL ASSETS

  CURRENT RATIO= ?


10. NET PROFIT= P 300,000, MARGIN= 0.1

   INCOME TAX EXPENSES= 25% OF OPERATING PROFIT

   OPERATING PROFIT MARGIN= ?


1
Expert's answer
2022-01-13T12:29:13-0500

Answer 6:

GROSS PROFIT MARGIN= 0.60

INVENTORY TURNOVER= 3

AVERAGE INVENTORY= P 450,000

NET PROFIT MARGIN= 0.12

"Inventory Turnover=\frac{Sales}{Average Inventories}"

"3=\frac{Sales}{450000}"

Sales = 1,350,000

"Net Profit=Sales\times Net Profit Margin=1350000\times0.12=162000"

 

Answer 7:

CURRENT RATIO = 4

DEBT RATIO= 0.2

EQUITY= P 600,000

NONCURRENT ASSETS= P 200,000

"Debt Ratio=\frac{Total Liabilities}{Total Assets}=0.2"

Equity = Total assets - Total liabilities

Total assets =Equity + Total liabilities

"\frac{Total Liabilities}{600000+Total Liabilities}=0.2"

"Total liabilities=120000+0.2Total liabilities"

"Total Liabilities=\frac{120000}{0.8}=150000"

"Total Assets=600000+150000=750000"

"Current Assets=750000-200000=550000"

"Current Liabilities=\frac{55000}{4}=137500"

"Noncurrent Liabilities=Total liabilities - Current Liabilities"

"Noncurrent Liabilities=150000-137500=12500"


Answer 8:

AVERAGE DAYS IN INVENTORY= 20

"Inventory Turnover=\frac{365}{20}=18.25"


Answer 9:

CURRENT LIABILITIES= 50% OF TOTAL ASSETS

CURRENT ASSETS= 75% OF TOTAL ASSETS

Let X = Total assets

Current assets = 0.75X

Current liabilities = 0.5X

"Current Ratio=\frac{0.75X}{0.5X}=1.5"


Answer 10:

NET PROFIT= P 300,000

MARGIN= 0.1

INCOME TAX EXPENSES= 25% OF OPERATING PROFIT

OPERATING PROFIT MARGIN= ?

"Revenue=\frac{300,000}{0.1}=3,000,000"

"Operating Profit=\frac{300,000}{1-0.25}=400,000"

"Operating Profit Margin=\frac{400,000}{3,000,000}=0.133\approx13.3\%"